OpenStudy (anonymous):

A small boat travels 12mph in still water. It takes 4hrs longer to travel 45 miles going upstream than downstream. Find the rate of the current. (Hint: 12+x= rate going down; 12-x= rate going up) round answer to nearest tenth. *I know you have to use the quadratic formula at some point but thats about it.

hero (hero):

Teacher sounds like someone who gives students partial answers on the test

OpenStudy (anonymous):

t=d/r The difference between the time up and the time down is 4 hours. Solve the following for c, the speed of the current: \[\frac{45}{12-c}-\frac{45}{12+c}= 4\]